A random sample is drawn from a population of known standard deviation 11.3. Construct 90%confidence interval for the population mean based on the information given (not all of the information given need be used).



a. n=36, x¯=105.2, s=11.2n=36, x¯=105.2, s=11.2




b. n=100, x¯=105.2, s=11.2n=100, x¯=105.2, s=11.2

1
Expert's answer
2022-05-03T13:31:18-0400

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

a) The critical value for "\alpha = 0.10, df=n-1=35" degrees of freedom is "t_c=z_{1\u2212\u03b1\/2;n\u22121}=1.689572."

"CI=(105.2-1.689572\times\dfrac{11.2}{\sqrt{36}},"

"105.2+1.689572\times\dfrac{11.2}{\sqrt{36}})"

"=(102.046,108.354)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "102.046< \mu < 108.354," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(102.046,108.354)."


b) The critical value for "\alpha = 0.10, df=n-1=99" degrees of freedom is "t_c=z_{1\u2212\u03b1\/2;n\u22121}= 1.660391."

"CI=(105.2- 1.660391\times\dfrac{11.2}{\sqrt{100}},"

"105.2+ 1.660391\times\dfrac{11.2}{\sqrt{100}})"

"=(103.340,107.060)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "103.340< \mu < 107.060," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(103.340,107.060)."